OpenStudy (anonymous):

Simplify: 5{3[2x-6]-3}

OpenStudy (anonymous):

Multiply 3(2x-6) using distributive property = 6x-18 Now you have 5(6x-18-3) so combine like terms. = 5(6x-21) Then use distributive property again = 30x-105